OpenStudy (leozap1):

*Fan & Medal to who ever helps* What is the simplified form of -3z^2 (z + 2) -4(z^2 + 1)? Please do NOT give me the answer explain to me how to get the answer my self

OpenStudy (leozap1):

What do I need to do first in this?

OpenStudy (anonymous):

in layman's terms.. break the bracket first..

OpenStudy (anonymous):

like \[-3z^2 (z+2) will be -3z^3 -6z^2\]

OpenStudy (anonymous):

try it your self... its easy..

OpenStudy (leozap1):

How did you get that?

OpenStudy (anonymous):

multiply (z+2) with -3z^2 1st step.. multiply z with z^2 you will get z^2+1 ie Z^3 since both base are same (ie z) so power will be added

OpenStudy (leozap1):

Sorry but I am still confused I worked it out and got -6z^3 - 4z^2 and that is not one of the answer choices

OpenStudy (anonymous):

ok Try again from basic you may be missing some thing while finding product

OpenStudy (leozap1):

When I multiply it out I get -6z^3 - 4z^2 I don't know what I am doing wrong

OpenStudy (anonymous):

write full how you got to this answer

OpenStudy (leozap1):

For the first part -3z^2 (z + 2) I multiplied it and got -3z^3 * 2 and got -6z^3 For the second part -4 (z^2 + 1) I multiplied it and got -4z^2

OpenStudy (anonymous):

no its wrong you have to multiply -3z^2 with both z and 2 like this \[-3z^2 \times z +-3z^2 \times 2= -3 z^3 +-6z^2\] which further simplified as \[-3 z^3 -6z^2\]

mathslover (mathslover):

Note this down : a(b+c) = ab + ac --> distributive property

OpenStudy (pokemon23):

Wouldn't you just make the problem more easier if you just multiply the parentheses ?

OpenStudy (pokemon23):

Yeah I'm right

mathslover (mathslover):

Now see, we will break the expression in 2 parts : \(-3z^2(z+2) \) + (\(-4(z^2+1)\) ) First solving for the first part : \(-3z^2(z+2)\)

OpenStudy (leozap1):

Ok what I got now is -3z^3-6z^2-4z^2-4

mathslover (mathslover):

Right! Good work.

mathslover (mathslover):

Now you just have to simplify it further.

OpenStudy (leozap1):

How can I do that?

mathslover (mathslover):

See, we have to simplify those terms which are like that is which have same power of the variable. Can you name those terms?

OpenStudy (leozap1):

I don't know sorry

mathslover (mathslover):

no problem, see : x^2 and 3x^2 are called like terms because they have same power of the variable (x) Similarly, in this expression : -3z^3-6z^2-4z^2-4 , we have two terms as the like terms : -6z^2 and -4z^2 as they both have same power of the variable z .

OpenStudy (leozap1):

Yah but I cant add the exponent because the bases are different

mathslover (mathslover):

yes but you can take the (z^2) as common : z^2(-6-4)

OpenStudy (leozap1):

So its -10z^2?

OpenStudy (leozap1):

So is the full answer -3z^3 - 10z^2 - 4?

mathslover (mathslover):

Correct

OpenStudy (leozap1):

Yay :) Thnxs for all the help

mathslover (mathslover):

You're very welcome :) , Good luck

OpenStudy (pokemon23):

Mission success

mathslover (mathslover):

Haha #MissionAnswerTheQuest. Accomplished SIR! :D

OpenStudy (anonymous):

really #Mission Success